OpenStudy (lgbasallote):

A gas evolved from the fermentation of glucose is found to effuse through a porous barrier in 15.0 min. Under the same conditions of temperature and pressure, it takes an equal volume of N2 12.0 min to effuse through the same barrier. Calculate the molar mass of the gas.

OpenStudy (lgbasallote):

what's given here is time not rate so i use the formula \[\huge \frac {t_1}{t_2} = \sqrt{\frac{M_1}{M_2}}?\] or is it \[\huge \frac{t_2}{t_1}= \sqrt{\frac{M_1}{M_2}}\] i cant ind the formula in the net..

OpenStudy (lgbasallote):

find*

OpenStudy (dominusscholae):

Assuming both gasses must diffuse the same measure of distance: The top one.

OpenStudy (lgbasallote):

ahh so time is directly proportional to molar mass huh

OpenStudy (dominusscholae):

^Square root of molar mass, yes, like in the first formula.

OpenStudy (lgbasallote):

\[M_1 = 43.78 g/mol\] is that right?

OpenStudy (lgbasallote):

@dominusscholae how would i find the root-mean-square of N2? do you have an idea how?

OpenStudy (dominusscholae):

You mean root-mean-square velocity?

OpenStudy (lgbasallote):

yup

OpenStudy (dominusscholae):

It's \[\sqrt{3RT/M}\] M the mass of N2

OpenStudy (lgbasallote):

is that a formula?`

OpenStudy (dominusscholae):

Yep

OpenStudy (lgbasallote):

i dont think that's applicable here though...i have no R or T right?

OpenStudy (dominusscholae):

You don't need it. You have the equation above to help you lol.

OpenStudy (dominusscholae):

In your first post.

OpenStudy (lgbasallote):

\[\frac{t_1}{t_2} = \sqrt{\frac{M_1}{M_2}}?\]

OpenStudy (lgbasallote):

how can i use that?

OpenStudy (dominusscholae):

You have the time of effusion for N2, the time of effusion for the unknown gas, and the mass of N2 (which you can calculate).

OpenStudy (lgbasallote):

i do not see where you are going..

OpenStudy (dominusscholae):

Plug t1 = 15.0 min , plug t2 = 12.0 min , plug M2 = 28.02 g/mol

OpenStudy (lgbasallote):

wouldnt M1 be the only thing left then?

OpenStudy (lgbasallote):

how would plugging it in help me find the root-mean-square speed?

OpenStudy (lgbasallote):

\[\frac{15}{12} = \sqrt{\frac{M_1}{28.02}}\] how can i solve for rms??

OpenStudy (mimi_x3):

\[\frac{15}{12} = \frac{\sqrt{M_{1}}}{\sqrt{28.02}} \] you should be able to solve for \(M_1\) now :)

OpenStudy (lgbasallote):

im not solving for M1 i've already solved for m1

OpenStudy (lgbasallote):

im asking for the root-mean-square

OpenStudy (mimi_x3):

define "root-mean-square"?

OpenStudy (lgbasallote):

idk...

OpenStudy (lgbasallote):

im asking arent i

OpenStudy (mimi_x3):

http://en.wikipedia.org/wiki/Root_mean_square i dont see how that is related to chemistry :/

OpenStudy (lgbasallote):

so you dont know how to solve this @Mimi_x3 huh...simple algebra she says -_-

OpenStudy (anonymous):

15/12 = 1.25 1.25 x sqrt 28.02 = 6.6 6.6^2= 43.78, more or less the molar mass of CO2. Mimi was right, just use algebra to solve for m1 but apparently you're incapable. idiot